find an equation of the line that contains the point (2,-3) and has the same slope as the line 3x-y+4=0

Answer by mamiya(56)


this is a general slope of a line: y - y' = m( x-x') where m is the slope of the line and (x', y') is a point of a line.
for this equation , we already know (x', y') , it is (2, -3)
now let's find m.
we know also that the slope of the line we are looking for is the same as the slope of 3x-y+4=0. so m=3 because the slope of 3x-y+4=0 is 3
thus, y-(-3) = 3(x-2) --> y+3 = 3x -6
--> y+3 - y-3 = 3x -6 -y -3
--> 0 = 3x -y -9
so the equation of that line is 3x - y - 9 =0